OpenStudy (anonymous):

log(2x-1) + log5 = 1

OpenStudy (nubeer):

log (5(2x-1)) =1 log has base 10.. take it on the other side \[5(2x-1) = 10^{1}\]

OpenStudy (anonymous):

Thanks! What about ln(x-1) + ln6 = ln(3x)

OpenStudy (nubeer):

when two logs are added they are being multiplied.. and when two logs are being subtracted they are divided.. in this one multiply 6 with x-1 making one ln on each side... whenever there is 1 ln on both sided u can cancel them

OpenStudy (anonymous):

so it's the same as log? [6(x-1)] = 3x ?

OpenStudy (nubeer):

yes same as log.. and yes it's right.

OpenStudy (anonymous):

the is 1/3 ?

OpenStudy (nubeer):

hmm no..6x-6=3x try to solve now..

OpenStudy (anonymous):

6x-6=3x, 3x = 6, x=6/3 x=2 that was a dumb moment! Thanks!!

OpenStudy (anonymous):

in f(x) =(x^3 - 3x^2 + 6) / (x+2) my answer is x^2-5x+10- (14/x+2) did I do this correctly?

OpenStudy (nubeer):

lol well u can try wolfram if you just want to check answers.. , but ya i think u did it right.

OpenStudy (anonymous):

Yeah, They give me the answer but sometimes I get the answer correct but not the steps. I'm just trying to be sure I have the procedure correct for finals on Thursday. Thanks for your help!!

OpenStudy (nubeer):

no problem at all. you are most welcome.